# Diffusion-based Adversarial Identity Manipulation for Facial Privacy Protection

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33th ACM International Conference on Multimedia; October 27–31, 2025; Ireland, Dublinisbn: 978-1-4503-XXXX-X/2018/06copyright: noneccs: Security and privacy Privacy protections
@@ 1. Introduction
Recent works have explored noise-based adversarial examples to protect facial privacy by introducing adversarial perturbations to source images to deceive FR systems (Rajabi et al., 2021; Szegedy et al., 2013a; Yang et al., 2021). However, these adversarial perturbations are typically constrained to $l_{p}$ norm in the pixel space, resulting in visible artifacts and poor image quality (Yang et al., 2021). Unlike noise-based methods, unrestricted adversarial examples (Hu et al., 2022; Shamshad et al., 2023; Yin et al., 2021; Zhu et al., 2019) are not confined by perturbation budgets, allowing for more natural modifications (Bhattad et al., 2019a; Song et al., 2018a; Xiao et al., 2018a).
@@
To address these challenges, we propose DiffAIM. Our approach leverages a pre-trained diffusion model (Ho et al., 2020; Song et al., 2020a) to generate natural and highly transferable adversarial faces, without requiring extra reference information or retraining. DiffAIM manipulates facial identity within the diffusion model’s latent space, leveraging its effective generative prior to produce more natural images. Specifically, we first employ edit-friendly latent mapping to map the source face image into a low-dimensional latent space. This step provides a good initialization for manipulation and the capability for faithful reconstruction. We then perform identity manipulation via iterative and gradient-based adversarial guidance injection during the reverse diffusion process. This iterative guidance allows for controllable and progressive steering toward the desired adversarial distribution. The guidance is derived from a carefully formulated objective for effective impersonation and visual naturalness. This objective promotes identity convergence towards a specific target, utilizing an adaptive ensemble of diverse FR models to enhance black-box transferability. Concurrently, it encourages semantic divergence from the source by utilizing intermediate features from the denoising U-Net (Ronneberger et al., 2015), facilitating natural-looking impersonation. We further incorporate structure-preserving regularization to preserve facial structure during manipulation. Finally, an identity-sensitive timestep truncation strategy is introduced to balance adversarial effectiveness and visual quality. In summary, our main contributions are:
@@ 2.1. Adversarial Attacks on Face Recognition
Adversarial attacks (Goodfellow et al., 2014; Szegedy et al., 2013b) present a key strategy for protecting facial privacy against unauthorized FR systems. These methods introduce carefully crafted perturbations to input images during inference to mislead the FR systems. Early methods often focused on noise-based attacks (Yang et al., 2021; Oh et al., 2017; Zhou et al., 2023; Zhong and Deng, 2022), which add perturbations typically constrained by $l_{p}$ norms. While demonstrating feasibility, methods like TIP-IM (Yang et al., 2021), which optimizes adversarial identity masks, often result in visually perceptible artifacts, limiting their practical applicability.
Another category of solutions employs unrestricted adversarial examples (Bhattad et al., 2019b; Kakizaki and Yoshida, 2019; Song et al., 2018b; Xiao et al., 2018b; Yuan et al., 2022; Zhao et al., 2020), which are not constrained by the perturbation norm in pixel space, thereby improving image quality (Bhattad et al., 2019b; Song et al., 2018b; Xiao et al., 2018b). Within this category, makeup-based methods attempt to embed adversarial perturbations into synthesized facial makeup. For instance, Adv-Makeup (Yin et al., 2021) applies adversarial patterns to local regions (e.g., eyeshadow). Other methods, such as AMT-GAN (Hu et al., 2022), CLIP2Protect (Shamshad et al., 2023), and DiffAM (Sun et al., 2024) utilize generative models guided by reference images, textual prompts, or learned makeup priors to create adversarial makeup. However, these techniques often require external user-specified makeup reference information and can generate unnatural or heavy makeup appearances. Furthermore, their typical focus on local attributes and dependence on makeup generation can limit their black-box transferability. In contrast, Adv-Diffusion (Liu et al., 2024) utilizes the inpainting capabilities (Wang et al., 2023; Couairon et al., 2022a) of diffusion models to generate higher-quality adversarial faces, but it exhibits limited attack transferability and destruction on the background content. GIFT (Li et al., 2024) employs global facial manipulation via GAN inversion (Kang et al., 2021) optimization. While achieving improved transferability, this approach often introduces substantial and unrealistic alterations to the original face. In this work, we propose DiffAIM to craft natural-looking adversarial faces without requiring extra reference information or model retraining.
@@ 2.2. Diffusion Methods
Diffusion models (Ho et al., 2020; Song et al., 2020a, b) are a class of probabilistic generative models renowned for their ability to synthesize high-quality images. The reverse process can be viewed as iterative denoising, transforming a simple prior distribution into a complex target distribution. To enhance computational efficiency, latent diffusion models (LDMs) (Rombach et al., 2022) perform the diffusion process within a compressed latent space learned by an autoencoder (Kingma et al., 2013; Sohn et al., 2015).
The powerful generative capabilities of diffusion models have led to their widespread adoption in various tasks, including text-to-image synthesis (Saharia et al., 2022; Chen et al., 2024c; Rombach et al., 2022), image editing (Couairon et al., 2022b; Avrahami et al., 2022; Kim et al., 2022) and image super-resolution (Gao et al., 2023; Li et al., 2022). Further, recent research integrates adversarial example generation with the diffusion process to produce high-quality adversarial examples (Chen et al., 2024a, b, 2023; Dai et al., 2024). For instance, Advdiff (Dai et al., 2024) uses a conditional diffusion model, adding adversarial perturbations during the denoising process starting from Gaussian noise. ACA (Chen et al., 2023) modifies the initial latent code before commencing the reverse diffusion process to embed adversarial semantics. Inspired by these works, we utilize a latent diffusion model and introduce a fine-grained iterative guidance mechanism to manipulate facial identity, aiming to generate natural yet effective adversarial faces.
@@ 3.1. Problem Formulation
Figure 2. Overview of the proposed DiffAIM. We first employ edit-friendly latent mapping to obtain the initial latent code and consistent noise maps for reconstruction. Next, adversarial guided diffusion is used to generate natural adversarial face images by injecting gradient-based guidance during the reverse diffusion process.
@@ 3.2. Edit-Friendly Diffusion Inversion
Diffusion inversion techniques (Dhariwal and Nichol, 2021; Mokady et al., 2023) are widely used in diffusion-based image editing tasks (Mokady et al., 2023; Couairon et al., 2022b). These techniques invert a real image into the latent space of a diffusion model, often by determining some initial noise that can be denoised to reconstruct the image. The noise can then be perturbed or modified during the generation process to edit the image. Recent work has shown that inverting the DDPM scheduler can achieve faithful reconstruction and few-step editing and propose the edit-friendly inversion (Huberman-Spiegelglas et al., 2024).
@@
where $\{z_{t}\}$ are iid standard normal vectors, $\sigma_{t}$ is derived from the noising schedule, and $\hat{\mu}_{t}(x_{t})$ is derived from the output of denoising network, commonly implemented as a U-Net (Ronneberger et al., 2015).
@@ 4.1. Overview
The subsequent sections detail the edit-friendly latent mapping and adversarial guidance injection process (Sec. 4.2), the objective function (Sec. 4.3), and the timestep truncation strategy (Sec. 4.4).
@@ 4.2. Adversarial Identity Manipulation
The update rule combines the standard reverse diffusion step (Eq. 2) using pre-computed $z_{t}$ with the adversarial guidance $\mathcal{G}_{t}$. The adversarial latent $\hat{x}_{t-1}$ for the previous timestep is computed as
@@ 4.3. Adversarial Identity Guidance Optimization
We now detail the formulation and optimization of the objective function used to compute the adversarial guidance gradient. The goal is to steer the generation towards effective impersonation while maintaining visual naturalness.
@@
Recent work (Kwon et al., 2022) has demonstrated that manipulating the deepest features within the intermediate U-Net block can lead to meaningful semantic edits without degrading image quality. Inspired by this, we hypothesize that leveraging high-level semantic features from U-Net for the divergence component can effectively push the generated image away from the source without causing significant visual degradation, thereby facilitating natural-looking impersonation. Consequently, we propose a refined adversarial loss that combines a FR-driven identity convergence term with a U-Net-driven semantic divergence term
@@
where $x_{t}$ is the benign inverted latent code produced in the latent mapping process, and $\hat{x}_{t}$ is the corresponding adversarial latent code, $\mathcal{U}_{\theta}^{\text{mid}}(\cdot)$ represents the features output from the intermediate block of the frozen U-Net. This formulation combines the explicit discriminative power of the FR model with the rich semantic and generative priors learned by U-Net, leading to effective adversarial results while better maintaining visual naturalness compared to the naive objective.
@@ 16      $\hat{x}_{t-1}=x_{t-1}+\mathcal{G}_{t}^{N_{a}}$;
Adaptive Ensemble Strategy. To further enhance the transferability of the generated adversarial faces against unseen black-box FR systems, we select $N_{m}$ pre-trained FR models $\{\mathcal{M}_{i}\}_{i=1}^{N_{m}}$ with high accuracy on public face datasets, serving as white-box models to simulate the decision boundaries of potential target models under the black-box setting. Specifically, the final form of adversarial loss is
@@
Structural Preservation Loss. Adversarial manipulation at each timestep can inadvertently distort the underlying facial structure, resulting in unnatural results. Inspired by findings (Tumanyan et al., 2022) that self-attention layers can capture structure information within images, we further introduce regularization to maintain structural consistency between the adversarial latent $\hat{x}_{t}$ and the benign latent $x_{t}$. We minimize the difference between their respective self-attention maps
@@ 4.4. Identity-Sensitive Timestep Truncation
We begin with the fact that diffusion models typically establish low-frequency global facial structures in earlier steps before refining high-frequency details concentrated in identity-sensitive regions (e.g., eyes, nose, cheek, Fig. 3 (a) and (b)), which cognitive psychology identifies as vital for identity discrimination. Second, we analyzed the attention patterns of the FR model across timestep using Grad-CAM (Selvaraju et al., 2017), as shown in Fig. 3 (c). This revealed that the model’s attention gradually shifts from a dispersed pattern to a more focused one, concentrating on these key identity-sensitive regions during an intermediate phase of the reverse diffusion process, approximately from $0.5T$ down to $0.1T$.
@@ 5.1. Experimental Setup
Datasets: We conduct experiments for both face verification and identification tasks. Face Verification: we use CelebA-HQ (Karras et al., 2017) and LADN (Gu et al., 2019) as our test sets. Specifically, we select a subset of 1000 images with different identities from CelebA-HQ and divide them into 4 groups, with each group of images impersonating a target identity provided by (Hu et al., 2022). Similarly, for LADN, we divide the 332 images into 4 groups for attack on different target identities. Face Identification: we use 500 randomly selected images with different identities from CelebA-HQ as the probe set. The gallery set is constructed using the corresponding 500 images of the same identities along with the 4 target identities provided by (Hu et al., 2022).
Baseline Methods: We compare the proposed approach with multiple noise-based and unrestricted-based facial privacy protection methods. Noise-based methods include PGD (Madry et al., 2017), and TIP-IM (Yang et al., 2021). Unrestricted-based methods include Adv-Makeup (Yin et al., 2021), AMT-GAN (Hu et al., 2022), CLIP2Protect (Shamshad et al., 2023), DiffAM (Sun et al., 2024), and GIFT (Li et al., 2024).
Target Models: We aim to protect facial privacy by attacking deep FR models in the block-box settings. Specifically, our study targets widely used FR models, including IRSE50 (Hu et al., 2018), IR152 (Deng et al., 2019), FaceNet (Schroff et al., 2015), and MobileFace (Chen et al., 2018). Meanwhile, we evaluate the privacy protection performance on commercial FR APIs, including Face++ and Aliyun.
@@
Implementation Details: We employ stable diffusion (Rombach et al., 2022) as the pre-trained diffusion model to generate adversarial face images. We utilize the DDIM sampler (Song et al., 2020a) with $T=100$ diffusion steps. We set the attack iterations $N_{a}=10$, the step size $\eta=3$, the truncated timestep $t_{s}=20$, and the hyper-parameter for structure preservation regularization is set to $\lambda=0.1$.
@@ 5.4. Ablation Study
Effect of Improved Adversarial Loss. As shown in Fig.5, the naive adversarial loss (Eq. 7) introduces perceptible artifacts, particularly around the eyes and mouth. Conversely, our proposed refined adversarial loss (Eq. 8) generates adversarial faces with a more natural appearance. Additionally, as shown in Tab. 4, integrating the U-Net-driven semantic divergence achieves enhanced adversarial effectiveness.
@@
Figure 8. Albation study for the truncated timestep $t_{s}$. Image quality metrics ($\dagger$) are scaled for visual trend comparison.
Selection of truncated timestep. We analyze the impact of the truncated timestep $t_{s}$. As shown in Fig. 8, injecting adversarial guidance early yields higher ASR but degrades image quality. Specifically, injecting guidance throughout the reverse diffusion process (i.e., $t_{s}=T$) introduces noticeable artifacts. Conversely, delaying guidance improves visual quality but causes a decline in ASR. The results reveal an optimal range around $t_{s}\in[0.2T,0.3T]$, where ASR remains high while achieving excellent visual quality. This validates the effectiveness of our identity-sensitive timestep truncation strategy.

### v3

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755478ccs: Security and privacy Privacy protections
@@ 1. Introduction
Recent works have explored noise-based adversarial examples (Zhou et al., 2025, 2024b) to protect facial privacy by introducing adversarial perturbations to source images to deceive FR systems (Rajabi et al., 2021; Szegedy et al., 2014; Yang et al., 2021). However, these adversarial perturbations are typically constrained to $l_{p}$ norm in the pixel space, resulting in visible artifacts and poor image quality. Unlike noise-based methods, unrestricted adversarial examples (Hu et al., 2022; Shamshad et al., 2023; Yin et al., 2021; Zhu et al., 2019) are not confined by perturbation budgets, allowing for more natural modifications (Bhattad et al., 2019; Song et al., 2018; Xiao et al., 2018).
@@
To address these challenges, we propose DiffAIM [^1] The code is available at https://github.com/Liqin-Wang/DiffAIM. Our approach leverages a pre-trained diffusion model (Ho et al., 2020; Song et al., 2021a) to generate natural and highly transferable adversarial faces, without requiring extra reference information or retraining. DiffAIM manipulates facial identity within the diffusion model’s latent space, leveraging its effective generative prior to produce more natural images. Specifically, we first employ edit-friendly latent mapping to map the source face image into a low-dimensional latent space. This step provides a good initialization for manipulation and the capability for faithful reconstruction. We then perform identity manipulation via iterative and gradient-based adversarial guidance injection during the reverse diffusion process. This iterative guidance allows for controllable and progressive steering toward the desired adversarial distribution. The guidance is derived from a carefully formulated objective for effective impersonation and visual naturalness. This objective promotes identity convergence towards a specific target, utilizing an adaptive ensemble of diverse FR models to enhance black-box transferability. Concurrently, it encourages semantic divergence from the source by utilizing intermediate features from the denoising U-Net (Ronneberger et al., 2015), facilitating natural-looking impersonation. We further incorporate structure-preserving regularization to preserve facial structure during manipulation. Finally, an identity-sensitive timestep truncation strategy is introduced to balance adversarial effectiveness and visual quality. In summary, our main contributions are:
@@ 2.1. Adversarial Attacks on Face Recognition
Adversarial attacks (Zhou et al., 2023b, a, 2024a) present a key strategy by introducing perturbations to input images during inference to mislead the FR systems. Early methods focused on noise-based attacks (Oh et al., 2017) constrained by $l_{p}$ norms, but often produce perceptible artifacts.
Another category of solutions employs unrestricted adversarial examples (Kakizaki and Yoshida, 2019), which are not constrained by the perturbation norm in pixel space, thereby improving image quality. Within this category, makeup-based methods (Yin et al., 2021; Hu et al., 2022; Shamshad et al., 2023) embed adversarial perturbations into synthesized facial makeup. However, these techniques often require external user-specified makeup reference information and can generate unnatural or heavy makeup appearances. Furthermore, their typical focus on local attributes and dependence on makeup generation can limit black-box transferability. In contrast, Adv-Diffusion (Liu et al., 2024) utilizes the inpainting capabilities (Wang et al., 2023) of diffusion models to generate higher-quality adversarial faces, but it exhibits limited attack transferability and destruction on the background content. GIFT (Li et al., 2024) employs global facial manipulation via GAN inversion (Kang et al., 2021) optimization, but it often introduces substantial and unrealistic alterations to the original face. In this work, we propose DiffAIM to craft natural-looking adversarial faces without requiring extra reference information or model retraining.
@@ 2.2. Diffusion Methods
Diffusion models (Ho et al., 2020; Song et al., 2021a, b) are a class of probabilistic generative models renowned for their ability to synthesize high-quality images. The reverse process can be viewed as iterative denoising, transforming a simple prior distribution into a complex target distribution. To enhance computational efficiency, latent diffusion models (LDMs) (Rombach et al., 2022) perform the diffusion process within a compressed latent space learned by an autoencoder (Kingma and Welling, 2014; Sohn et al., 2015).
The powerful generative capabilities of diffusion models have led to their widespread adoption in various tasks, including text-to-image synthesis (Saharia et al., 2022; Chen et al., 2024; Rombach et al., 2022), image editing (Couairon et al., 2023; Avrahami et al., 2022; Kim et al., 2022) and image super-resolution (Gao et al., 2023; Li et al., 2022). Further, recent research integrates adversarial example generation with the diffusion process to produce high-quality adversarial examples (Chen et al., 2025). For instance, Advdiff (Dai et al., 2024) uses a conditional diffusion model, adding adversarial perturbations during the denoising process starting from Gaussian noise. ACA (Chen et al., 2023) modifies the initial latent code before commencing the reverse diffusion process to embed adversarial semantics. Inspired by these works, we utilize a latent diffusion model and introduce a fine-grained iterative guidance mechanism to manipulate facial identity, aiming to generate natural yet effective adversarial faces.
@@ 3.1. Problem Formulation
Figure 2. Overview of the proposed DiffAIM. We first employ edit-friendly latent mapping to obtain the initial latent code and consistent noise maps for reconstruction. Next, adversarial guided diffusion is used to generate natural adversarial face images.
@@ 3.2. Edit-Friendly Diffusion Inversion
Diffusion inversion techniques (Dhariwal and Nichol, 2021; Mokady et al., 2023) are widely used in diffusion-based image editing tasks. These techniques invert a real image into the latent space of a diffusion model, often by determining some initial noise that can be denoised to reconstruct the image. The noise can then be perturbed or modified during the generation process to edit the image. Recent work has shown that inverting the DDPM scheduler can achieve faithful reconstruction and few-step editing and propose the edit-friendly inversion (Huberman-Spiegelglas et al., 2024).
@@
where $\{z_{t}\}$ are iid standard normal vectors, $\sigma_{t}$ is derived from the noising schedule, and $\hat{\mu}_{t}(x_{t})$ is derived from the output of denoising network, commonly implemented as a U-Net.
@@ 4.2. Adversarial Identity Manipulation
The update rule combines the standard reverse diffusion step using pre-computed $z_{t}$ with the adversarial guidance $\mathcal{G}_{t}$. The adversarial latent $\hat{x}_{t-1}$ for the previous timestep is computed as
@@ 4.3. Adversarial Identity Guidance Optimization
We now detail the formulation and optimization of the objective function used to derive the adversarial guidance, aiming to steer the generation towards natural impersonation.
@@
Recent work (Kwon et al., 2023) has demonstrated that manipulating the deepest features within the intermediate U-Net block can lead to meaningful semantic edits without degrading image quality. Inspired by this, we hypothesize that leveraging high-level semantic features from U-Net for the divergence component can effectively push the generated image away from the source without causing significant visual degradation, thereby facilitating natural-looking impersonation. Consequently, we propose a refined adversarial loss that combines a FR-driven identity convergence term with a U-Net-driven semantic divergence term
@@
where $x_{t}$ is the benign inverted latent code produced in the latent mapping process, and $\hat{x}_{t}$ is the corresponding adversarial latent code, $\mathcal{U}_{\theta}^{\text{mid}}(\cdot)$ represents the features output from the intermediate block of the U-Net. This formulation combines the explicit discriminative power of FR models with the rich semantic and generative priors learned by U-Net, leading to more effective adversarial results and better visual naturalness compared to the naive ones.
@@ 16 $\hat{x}_{t-1}=x_{t-1}+\mathcal{G}_{t}^{N_{a}}$;
Adaptive Ensemble Strategy. To further enhance the transferability of the generated adversarial faces against unseen black-box FR systems, we select $N_{m}$ pre-trained FR models $\{\mathcal{M}_{i}\}_{i=1}^{N_{m}}$ with high accuracy on public face datasets, serving as white-box models to simulate the decision boundaries of potential target models under the black-box setting. The final form of adversarial loss is
@@
Structural Preservation Loss. Adversarial manipulation at each timestep can inadvertently distort the underlying facial structure. Inspired by findings (Tumanyan et al., 2022) that self-attention layers can capture structure information within images, we further introduce regularization to maintain structural consistency between the adversarial latent $\hat{x}_{t}$ and the benign latent $x_{t}$. We minimize the difference between their respective self-attention maps as
@@ 4.4. Identity-Sensitive Timestep Truncation
[…]
@@ 5.1. Experimental Setup
Implementation Details: We employ stable diffusion (Rombach et al., 2022) as the pre-trained diffusion model to generate adversarial face images. We utilize the DDIM sampler (Song et al., 2021a) with $T=100$ diffusion steps. We set the attack iterations $N_{a}=10$, the step size $\eta=3$, the truncated timestep $t_{s}=20$, and the hyper-parameter for structure preservation regularization is set to $\lambda=0.1$.
Datasets: We conduct experiments for both face verification and identification tasks. Face Verification: we use CelebA-HQ (Karras et al., 2018) and LADN (Gu et al., 2019) as our test sets. Specifically, we select a subset of 1000 images with different identities from CelebA-HQ and divide them into 4 groups, with each group of images impersonating a target identity provided by (Hu et al., 2022). Similarly, for LADN, we divide the 332 images into 4 groups for attack on different target identities. Face Identification: we use 500 randomly selected images with different identities from CelebA-HQ as the probe set. The gallery set is constructed using the corresponding 500 images of the same identities along with the 4 target identities provided by (Hu et al., 2022).
Baseline Methods: We compare the proposed approach with multiple noise-based and unrestricted-based facial privacy protection methods. Noise-based methods include PGD (Madry et al., 2018), and TIP-IM (Yang et al., 2021). Unrestricted-based methods include Adv-Makeup (Yin et al., 2021), AMT-GAN (Hu et al., 2022), CLIP2Protect (Shamshad et al., 2023), DiffAM (Sun et al., 2024), and GIFT (Li et al., 2024).
Target Models: We aim to protect facial privacy by attacking deep FR models in the block-box settings. Specifically, our study targets widely used FR models, including IRSE50 (Hu et al., 2018), IR152 (Deng et al., 2019), FaceNet (Schroff et al., 2015), and MobileFace (Chen et al., 2018). Meanwhile, we evaluate the performance on commercial FR APIs, including Face++ and Aliyun.
@@ 5.4. Ablation Study
Effect of Improved Adversarial Loss. As shown in Fig.5, the naive adversarial loss introduces perceptible artifacts, particularly around the eyes and mouth. Conversely, our proposed refined adversarial loss generates adversarial faces with a more natural appearance. Additionally, as shown in Tab. 4, integrating the U-Net-driven semantic divergence achieves enhanced adversarial effectiveness.
@@
Figure 8. Ablation study for the truncated timestep $t_{s}$. Image quality metrics ($\dagger$) are scaled for visual trend comparison.
Selection of truncated timestep. We analyze the impact of the truncated timestep $t_{s}$. As shown in Fig. 8, injecting adversarial guidance early yields higher ASR but degrades image quality. Specifically, injecting guidance throughout the reverse diffusion process (i.e., $t_{s}=T$) introduces noticeable artifacts. Conversely, delaying guidance improves visual quality but causes a decline in ASR. The results reveal an optimal range around $t_{s}\in[0.2T,0.3T]$, where ASR remains high while achieving excellent visual quality.
@@ 6. Conclusion
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (No.62441237, No.62172435).